OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 All Real Numbers No Solution 0 3

OpenStudy (solomonzelman):

2(x + 1) = 2x + 5 2x + 2 = 2x + 5 subtract 2x from both sides, and tell me what you think.

OpenStudy (anonymous):

its either no solution or 3 but i think its no solution

OpenStudy (solomonzelman):

You are thinking correctly, there is no solution.

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

You welcome !

OpenStudy (anonymous):

The table shows the solution to the equation |2x - 3| - 1 = 2: Step 1 |2x - 3| = 2 + 1 Step 2 |2x - 3| = 3 Step 3 2x - 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct

OpenStudy (solomonzelman):

What do you think ?

OpenStudy (anonymous):

can you help me with this one too @SolomonZelman

OpenStudy (solomonzelman):

yes, yes.. tel me, WHAT DO YOU THINK is the INCORRECT step ?

OpenStudy (anonymous):

step 3?

OpenStudy (solomonzelman):

BINGO !!! Right again !

OpenStudy (anonymous):

haha thanks

OpenStudy (solomonzelman):

You welcome, you are a smart girl :)